OpenStudy (anonymous):

Select the algebraic equation that correctly represents the given sentence and solve the equation . If 4 is subtracted from twice a number,the result is 12 less than the number

OpenStudy (anonymous):

\[x=number\]

OpenStudy (anonymous):

\[4-x+2=12-x\]

OpenStudy (nurali):

Let the number be x. Twice the number = 2x. "4 is subtracted from twice a number" => "4 is subtracted from 2x That means, 2x - 4 As the result is 12 less than number, that is : x - 12. Therefore , 2x - 4 = x-12

OpenStudy (anonymous):

this is the equation

OpenStudy (anonymous):

thanks